OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5

sam (.sam.):

2x + 3y = 4 2x = 4- 3y x = (4- 3y)/2 x - 2y = -5 (4- 3y)/2 - 2y = -5 y=2 x=-1

OpenStudy (anonymous):

2x + 3y = 4 x - 2y = -5 -> ( -1, 2)